# Residual Neural Terminal Constraint for MPC-based Collision Avoidance in Dynamic Environments

### v1

@@ Appendix D Simulation Experiments
The simulation experiments are performed in the Gazebo Harmonic 3D simulator. The task for the robot is to move from the initial state to the goal state while avoiding collisions with moving obstacles. We generate 100 different scenarios with random initialization of positions and velocities for six moving obstacles within the corridor that the robot needs to pass through. The obstacles move with constant speed and bounce back when they reach the edge of the corridor, so that all obstacles stay within the corridor all the time. The testing environment is shown in Fig. 9 and the motion executed by the robot for all MPC types with the prediction horizon of 10 steps (1 s) is visualized in Fig. 10.
The MPC local planner is implemented using the CasADi [29] framework within a ROS 2 package. To improve the real-time performance of the planner, we utilize CasADi’s feature to autogenerate C code and compile it before deployment, and additionally, we use the high-performance linear solver ma57 [30], which significantly improves computational time. The hypernetwork is implemented as a PyTorch model, the same as during the training phase, and achieves an inference speed of $\sim$2 ms on an RTX3090 GPU. However, integrating an MLP model with dynamically updated parameters as the MPC terminal constraint function is not trivial. Our solution is to implement the main network as a custom Python class using the operations provided by CasADi. In this way, the main network is integrated into the computational graph of the optimization problem, just like other functions. Also, the parameters of the main network are declared as parameters of the NLP solver, so they can be changed at every time step (same as the estimated and reference states).
@@
Figure 10: Visualization of the motion executed by the robot in all trials for different MPC types with the prediction horizon of 10 steps (1 s).
@@ Appendix E Hardware Experiments
The hardware experiments are performed using a Continental mobile robot equipped with a 3D LiDAR and stereo cameras. Complete computation and data processing, such as object recognition and tracking, robot localization, MPC optimization, etc., are performed using the onboard computer [^3] Intel NUC11PHi7 (CPU: 11th Gen Intel® Core™ i7-1165G7 @ 2.80GHz × 8; RAM: 16GB; GPU: NVIDIA GeForce RTX 2060).. As in the simulation experiments, the MPC planner is implemented using CasADi and deployed as a ROS 2 package. Also, the hypernetwork is deployed as a PyTorch model (with inference speed of $\sim$4.5 ms on an RTX2060 GPU), while the main network is implemented using CasADi functions. The discretization time of the MPC is $\delta t=0.1$ s and the control frequency is 20 HZ. Motion prediction of moving agents is performed using the constant velocity model.
During the hardware experiments, we conduct 10 trials for each method — 5 experiments with one pedestrian intersecting robot motion and 5 experiments with two pedestrians. The initial positions and velocities of pedestrians are random. The main difficulty of collision avoidance in those experiments arises from the limited actuation of the robot (relatively low maximal speed compared to moving obstacles), short prediction horizon, and complete onboard processing. In Fig. 11 we illustrate one example of a successfully avoided collision with a human, even though the MPC operates with a relatively short prediction horizon of 10 steps (1 s). To assess the quality of the executed trajectory, we visualize the motion of the robot for different baselines in Fig. 12. The results demonstrate significantly lower lateral deviation in the case of the proposed RNTC-MPC method compared to DCBF-MPC and VO-MPC, while SDF-MPC fails to avoid collision. An additional visualization is presented in Fig. 13.
Figure 11: Illustration of a successful collision avoidance of the proposed RNTC-MPC method with the prediction horizon of 10 steps (1 s) from the hardware experiments.

### v2

@@ Appendix D Simulation Experiments
The simulation experiments are performed using the Gazebo 3D simulator. The task for the robot is to move from the initial state to the goal state while avoiding collisions with moving obstacles. We generate 100 different scenarios with random initialization of positions and velocities for six moving obstacles within the corridor that the robot needs to pass through. The obstacles move with constant speed and bounce back when they reach the edge of the corridor, so that all obstacles stay within the corridor all the time. The testing environment is shown in Fig. 9 and the motion executed by the robot for all MPC types with the prediction horizon of 10 steps (1 s) is visualized in Fig. 10.
The MPC local planner is implemented using the CasADi framework [28] as a ROS 2 package with the control frequency of 20 Hz and sampling time $\delta t=0.1$ s. To improve the real-time performance of the planner, we utilize CasADi’s feature to autogenerate C code and compile it before deployment, and additionally, we use the high-performance linear solver MA57 [29], which significantly improves computational time. The hypernetwork is implemented as a PyTorch model, the same as during the training phase, and achieves an inference speed of $\sim$2 ms on an RTX3090 GPU. However, integrating an MLP model with dynamically updated parameters as the MPC terminal constraint function is not trivial. Our solution is to implement the main network as a custom Python class using the operations provided by CasADi. In this way, the main network is integrated into the computational graph of the optimization problem, just like other functions. Also, the parameters of the main network are declared as parameters of the NLP solver, so they can be changed at every time step (same as the estimated and reference states).
@@
Figure 10: Visualization of the motion executed by the robot during simulation experiments in all trials for different MPC types with the prediction horizon of 10 steps (1 s).
@@ Appendix E Hardware Experiments
The hardware experiments are performed using a Continental mobile delivery robot equipped with a 3D LiDAR and stereo cameras. Complete computation and data processing, such as object recognition and tracking, robot localization, MPC optimization, etc., are performed using the onboard computer [^4] Intel NUC11PHi7 (CPU: 11th Gen Intel® Core™ i7-1165G7 @ 2.80GHz × 8; RAM: 16GB; GPU: NVIDIA GeForce RTX 2060).. As in the simulation experiments, the MPC planner is implemented using CasADi and deployed as a ROS 2 package. Also, the hypernetwork is deployed as a PyTorch model (with inference speed of $\sim$4.5 ms on an RTX2060 GPU), while the main network is implemented using CasADi functions. The discretization time of the MPC is $\delta t=0.1$ s and the control frequency is 20 Hz. Motion prediction of moving agents is performed using the constant velocity model.
During the hardware experiments, we conduct 10 trials for each method — 5 experiments with one pedestrian intersecting robot motion and 5 experiments with two pedestrians. The initial positions and velocities of pedestrians are random. The main difficulty of collision avoidance in those experiments arises from the limited actuation of the robot (relatively low maximal speed compared to moving obstacles), short prediction horizon, and complete onboard processing. In Fig. 11 we illustrate one example of a successfully avoided collision with a human, even though the MPC operates with a relatively short prediction horizon of 10 steps (1 s). To assess the quality of the executed trajectory, we visualize the motion of the robot for different baselines in Fig. 12. The results demonstrate significantly lower lateral deviation in the case of the proposed RNTC-MPC method compared to DCBF-MPC and VO-MPC, while SDF-MPC fails to avoid collision. An additional visualization of the robot’s perception and estimated terminal set is presented in Fig. 13.
Figure 11: Illustration of a successful collision avoidance of the proposed RNTC-MPC method with the prediction horizon of 10 steps (1 s) during the hardware experiments.